OpenStudy (anonymous):

SOMEONE PLEASE PLEASE PLEASE HELP!! Will fan and medal help!! Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. (4 points each.) 1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 2. 12 + 42 + 72 + ... + (3n - 2)2 = For the given statement Pn, write the statements P1, Pk, and Pk+1. (2 points) 3. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (amistre64):

i believe your copy/paste is missing information

OpenStudy (anonymous):

Where?

OpenStudy (amistre64):

1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 2. 12 + 42 + 72 + ... + (3n - 2)2 = what does it equal? and the paste ate all of the ^s

OpenStudy (anonymous):

That equal sign is to represent you starting to work out the equation

OpenStudy (amistre64):

you cant prove the statements since they are incomplete

OpenStudy (anonymous):

How are they incomplete?

OpenStudy (amistre64):

prove that 3x =

OpenStudy (amistre64):

prove that the color of the train is .........

OpenStudy (amistre64):

prove that a tree is made of ........

OpenStudy (amistre64):

there should be formulas attached to the right side of the equal sign in order to make a comparison with and to be able to reference to when formating the induction process

OpenStudy (anonymous):

okay for number one it is equal to 4(4n+1)(8n+7)/6

OpenStudy (amistre64):

well, lets prove that for n=1 :)

OpenStudy (anonymous):

for number two it is equal to n(6n^2-3n-1)/2

OpenStudy (anonymous):

Okay how would we do that?

OpenStudy (amistre64):

replace n with a 1 .... does: 4(1) (4(1)+2) = 4(4(1)+1)(8(1)+7)/6 ??

OpenStudy (anonymous):

No.

OpenStudy (amistre64):

then as long as you posted it correctly, then we have proven that this is not true for ALL positive integers since we know it is not true for at least one of them

OpenStudy (anonymous):

Oh okay. So how would we solve number 2?

OpenStudy (amistre64):

same first step; let n=1, is the setup true for n=1?

OpenStudy (amistre64):

... + (3n - 2)2 = n(6n^2-3n-1)/2 ; let n=1 does: (3(1) - 2)2 = 1(6(1)^2 -3(1)-1)/2 ? i believe thats spose to be a ^2 onthe left ....

OpenStudy (anonymous):

Yes it is sorry

OpenStudy (amistre64):

then we have a true condition: 1 = 1 for n=1 assume it is true for some n=k; and see if we can format it the same way for the induction step

OpenStudy (anonymous):

How did you get 1=1 for n=1?

OpenStudy (amistre64):

i just work the math .... (3(1) - 2)^2 = 1(6(1)^2-3(1)-1)/2 (3 - 2)^2 = (6(1)-3-1)/2 (1)^2 = (6-4)/2 1 = 2/2 = 1 12 + 42 + 72 + ... + (3k - 2)^2 = k(6k^2-3k-1)/2

OpenStudy (anonymous):

So how do we solve it now with k?

OpenStudy (amistre64):

k is just a generic placeholder for "some value of n". If we can show that this same format is true for n=k+1, then we know that since its true for n=k=1, then its true for n=2, n=3, n=4, ...

OpenStudy (amistre64):

putting it into the k format seems a little redundant, but its just for generality aspects

OpenStudy (anonymous):

How would you write out the steps?

OpenStudy (amistre64):

like i did, assume its true for some n=k 12 + 42 + 72 + ... + (3k - 2)^2 = k(6k^2-3k-1)/2

OpenStudy (anonymous):

What would you do next?

OpenStudy (amistre64):

let P(k) equal that left side for a clean up P(k) = k(6k^2-3k-1)/2 now add the (k+1) element to each side P(k) + (3(k+1) - 2)^2 = k(6k^2-3k-1)/2 + (3(k+1) - 2)^2 the left side is equal to P(k+1); can we form the right side of it into a k+1 version as well? remember, all we know is a format; so we have to show that the format is consistent

OpenStudy (amistre64):

\[ \frac12k(6k^2-3k-1) + (3(k+1) - 2)^2\] \[ \frac12(6k^3-3k^2-k) + 9(k+1)^2 +4-12(k+1)\] \[ \frac12(6k^3-3k^2-k+ 18(k+1)^2 +8-24(k+1)) \] work the math .....

OpenStudy (amistre64):

typoed it .... we want to try to get it into the format: \[\frac12(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

Okay I get it now. How would you do number 3?

OpenStudy (amistre64):

its the same process, let n=1, let n=k, then proof it out for n=k+1

OpenStudy (anonymous):

So for n=1 it you would end up having 2(1)=1(1+1)?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

Okay. Thank you so much.

OpenStudy (amistre64):

youre welcome